# envname-P envname#1

### v1

@@ Abstract
Malicious behavior is often hidden in small, easily overlooked code fragments, especially within large and complex codebases. The cross-file dependencies of these fragments make it difficult for even powerful large language models (LLMs) to detect them reliably. We propose a graph-centric attention acquisition pipeline that enhances LLMs’ ability to localize malicious behavior. The approach parses a project into a code graph, uses an LLM to encode nodes with semantic and structural signals, and trains a Graph Neural Network (GNN) under sparse supervision. The GNN performs an initial detection, and through backtracking of its predictions, identifies key code sections that are most likely to contain malicious behavior. These influential regions are then used to guide the LLM’s attention for in-depth analysis. This strategy significantly reduces interference from irrelevant context while maintaining low annotation costs. Extensive experiments show that the method consistently outperforms existing methods on multiple public and self-built datasets, highlighting its potential for practical deployment in software security scenarios.
@@ 3.2 Attention Based Detection
$\mathcal{L}_{\text{pred}}=\log\left(1-p_{j,(1)}\right),$ (5)
where $\mathcal{L}_{\text{pred}}$ encourages the mask to only preserve the substructure that could lead the $G_{j}^{\text{code}}$ to be classified as malicious. To encourage sparsity and determinacy, we add size regularizers on the edge and feature masks:
@@
$\text{Att}(G^{\text{code}}_{j})=\Big(\big\{v\in\mathcal{V}_{j}^{\text{code}}|a_{v}<\gamma^{\text{node}}\big\},\big\{(v,w)\in\mathcal{E}_{j}^{\text{code}}|a_{v,w}<\gamma^{\text{edge}}\big\}\Big).$ (11)
@@ 4.2 Malicious Code Detection Capability Comparisons (RQ1, RQ2)
We first compared the accuracy of malicious code detection. Tables 1 and 2 present the detailed results. As observed, our method demonstrates superior performance on both public datasets and our own constructed dataset, outperforming baseline methods in the majority of scenarios and metrics. Notably, the improvement of our method over LLMs is significant, especially in the Precision metric, highlighting the ability of GMLLM to filter out redundant benign code interference, thereby validating the effectiveness of GMLLM. Furthermore, the comparison with other tools also indicates that, based on GMLLM, the LLM-based approach can effectively handle malicious code detection, particularly in the domain of Python code.
@@ 4.4 LLM Resource Consumption (RQ4)
We compared the resource consumption of GMLLM against other LLMs, using the number of tokens consumed as the primary metric. The results are presented in Table 4. As the data shows, GMLLM demonstrates the lowest resource usage in all scenarios, and the margin of improvement is significant. In some cases, GMLLM achieves a reduction in computational cost by several orders of magnitude, which clearly demonstrates the superiority of our method. Beyond token usage, we also profile the cost of the GNN-Explainer step itself. On MalCP, optimizing the masks for a single project takes on average 10–11 seconds (see Appendix E.1 for full statistics).

### v2

@@ Abstract
Large Language Models (LLMs) have significantly advanced code analysis tasks, yet they struggle to detect malicious behaviors fragmented across files, whose intricate dependencies easily get lost in the vast amount of benign code. We therefore propose a graph-centric attention acquisition pipeline that enhances LLMs’ ability to localize malicious behavior. The approach parses a project into a code graph, uses an LLM to encode nodes with semantic and structural signals, and trains a Graph Neural Network (GNN) under sparse supervision. The GNN performs an initial detection, and by interpreting these predictions, identifies key code sections that are most likely to contain malicious behavior. These influential regions are then used to guide the LLM’s attention for in-depth analysis. This strategy significantly reduces interference from irrelevant context while maintaining low annotation costs. Extensive experiments show that the method consistently outperforms existing approaches on multiple public and custom datasets, highlighting its potential for practical deployment in software security scenarios.
@@ 3.2 Attention Based Detection
$\mathcal{L}_{\text{pred}}=-\log\left(p_{j,(1)}+\epsilon\right),$ (5)
where minimizing $\mathcal{L}_{\text{pred}}$ encourages the mask to preserve the substructure that maximize $p_{j,(1)}$
To encourage sparsity and determinacy, we add size regularizers on the edge and feature masks:
@@
$\begin{split}\text{Att}(G^{\text{code}}_{j})=\Big(&\big\{v\in\mathcal{V}_{j}^{\text{code}}|a_{v}>\gamma^{\text{node}}\big\},\\ &\big\{(v,w)\in\mathcal{E}_{j}^{\text{code}}|a_{v,w}>\gamma^{\text{edge}}\big\}\Big).\end{split}$ (11)
@@ 4.2 Malicious Code Detection Capability Comparisons (RQ1, RQ2)
We first compared the accuracy of malicious code detection. Tables 1 and 2 present the detailed results. As observed, our method demonstrates superior performance on both public datasets and our own constructed dataset, outperforming baseline methods in the majority of scenarios and metrics. Notably, the improvement of our method over LLMs is significant, especially in the Precision metric, highlighting the ability of GMLLM to filter out redundant benign code interference, thereby validating the effectiveness of GMLLM. Furthermore, the comparison with other tools also indicates that, based on GMLLM, the LLM-based approach can effectively handle malicious code detection, particularly in the domain of Python code. Interestingly, on the Small dataset where baseline LLMs achieve near-perfect recall, GMLLM maintains comparable performance, indicating that the graph-guided mechanism does not introduce unnecessary overhead when code complexity is low. This absence of regression validates the generality of our approach across varying project scales.
@@ 4.4 LLM Resource Consumption (RQ4)
We compared the resource consumption of GMLLM against other LLMs, using the number of tokens consumed as the primary metric. Calculated via model-specific tokenizers, this metric strictly accounts for system prompts and formatting overheads (see Appendix D.3). The results are presented in Table 4. As the data shows, GMLLM demonstrates the lowest resource usage in all scenarios, and the margin of improvement is significant. In some cases, GMLLM achieves a reduction in computational cost by several orders of magnitude, which clearly demonstrates the superiority of our method. Beyond token usage, we also profile the cost of the GNN-Explainer step itself. On MalCP, optimizing the masks for a single project takes on average 10–11 seconds (see Appendix E.1 for full statistics).
@@ 4.6 Deeper Insight
This pattern also aligns with our observation that GMLLM does not regress on Small dataset. When code complexity is low, informative structures tend to be denser and more coherent, so graph guidance avoids over-filtering and preserves strong performance, supporting the generality of our approach.
@@ D.2.4 Runtime Environment and Reproducibility
D.3 Token Calculation Details
To ensure reproducibility and accurate cost estimation, token counts were computed using model-specific pre-trained tokenizers, avoiding character-based heuristics. For OpenAI models (e.g., GPT-4o), we used tiktoken with the o200k_base encoding; for open-source models (e.g., Llama-3), we employed the official Hugging Face tokenizers. This accounts for model-specific vocabulary as well as chat-template-mandated control tokens (e.g., <|im_start|>).
The reported token counts reflect the full input context of a single-turn inference pass, including the system prompt, the user query (containing the extracted subgraph), and template-induced formatting tokens. No chunking or context accumulation was applied; each prediction is based on a single, holistic prompt to assess the model’s capacity to reason over the compressed graph representation.